OpenStudy (anonymous):

Which of the point-slope equations below is correct for the line that passes through points (-4,-4) and (1,3)? Check all that apply.

OpenStudy (anonymous):

A. y - 3 = 7/5(x - 1) B. y - (-4) = 7/5(x - (-4)) C. y - 3 = 7/5(x + 1) D. y + 4 = 7/5(x - 4) E. y + 4 = 7/5(x + 4) F. y + 4 = 5/7(x + 4)

OpenStudy (anonymous):

not that your slope m is (Y2-y1)/(x2-x1) in your case (3+4)/(1+4) =7/5 now just pick a point and plug it into your point slope formula.

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

i just need to know the answer but thanks

OpenStudy (anonymous):

a and e, you'll never learn if you don't do it yourself though

OpenStudy (anonymous):

i know